# SCALA: Sparsification-based Contrastive Learning for Anomaly Detection on Attributed Networks

### 03 Jan 2024 v1

@@ Abstract
Anomaly detection on attributed networks aims to find the nodes whose behaviors are significantly different from other majority nodes. Generally, network data contains information about relationships between entities, and the anomaly is usually embodied in these relationships. Therefore, how to comprehensively model complex interaction patterns in networks is still a major focus. It can be observed that anomalies in networks violate the homophily assumption. However, most existing studies only considered this phenomenon obliquely rather than explicitly. Besides, the node representation of normal entities can be perturbed easily by the noise relationships introduced by anomalous nodes. To address the above issues, we present a novel contrastive learning framework for anomaly detection on attributed networks, SCALA, aiming to improve the embedding quality of the network and provide a new measurement of qualifying the anomaly score for each node by introducing sparsification into the conventional method. Extensive experiments are conducted on five benchmark real-world datasets and the results show that SCALA consistently outperforms all baseline methods significantly.
@@ 1 Introduction
Tasks on networks have drawn much attention in recent years Tan et al. (2019), particularly for anomaly detection on the network because there are many complex interaction patterns between entities hard to process Kim et al. (2022). Unlike in other domains, the anomaly detection in networks requires the identification of multiple anomaly types, including attribute anomaly and structure anomaly Ding et al. (2019). The former describes the node with the unusual condition of attributes and the latter means the node that has the abnormal situation of the network topology. How to simultaneously identify these anomalies remains a major current focus.
@@
Thanks to the increasing development of GNN which provides a tool to extract the topology and attribute knowledge at the same time Wu et al. (2021). Many researchers introduced GNN into anomaly detection on attributed networks tasksKim et al. (2022). In general, there are three main categories of methods. First, the Auto-Encoder based method assumes that the attribute and the topology pattern of a normal node can be correctly reconstructed by the decoder while the abnormal nodes cannotDing et al. (2019). Second, the One-Class SVM based method believes that all the normal nodes can be mapped in the boundary of a hypersphere with the appropriate loss function guidance, while the abnormal node will be expelled out of the boundaryWang et al. (2021). At last, the Contrastive Learning based (CL-based) method determines the degree of node anomaly by calculating the agreement score between the target node and its surrounding neighbor nodes. For normal nodes, their dense embeddings should maintain a matching condition with their positive subgraph embeddings, while abnormal nodes do not retain such a conditionLiu et al. (2021). A number of studies have shown that the CL-based method is currently the best framework of all because it can comprehensively and intuitively leverage the matching relationship between the target node and its neighborsLiu et al. (2021); Jin et al. (2021); Zheng et al. (2021); Zhang et al. (2022); Duan et al. (2023a).
@@
It is generally accepted that the most nodes of the real-world network satisfy homophily assumption, which means that connected nodes tend to have a high degree of similarity McPherson et al. (2001). However, as shown in Figure 1(a), the abnormal nodes don’t. Surprisingly, using this discipline as the criteria for spotting anomalies has remained underexplored in previous studies. Besides, as illustrated in Figure 1(b), the noise information of abnormal nodes will cause divergence when obtaining the embedding because most of the CL-based methods inevitably pick the abnormal node during sampling the contrastive pair. Therefore, the agreement calculation will have an under-best performance.
@@
  • We creatively introduce a new way to identify anomalies which complement the structural information neglected by the conventional CL-based method.
  • Extensive experiments have been conducted on five real-world datasets, whose results show that SCALA consistently outperforms all baseline methods significantly.
@@ 2.1 Anomaly Detection on Attributed Network
With the rapid development of deep learning, researchers also tried to adopt deep learning techniques to detect anomalies in networks. For instance, Ding et al. (2019) adopts an autoencoder to spot anomalies by measuring the reconstruction errors of nodes from both the structure and the attribute perspectives. Yuan et al. (2021) explicitly introduces the high-order motif structure information into the autoencoder scheme. Luo et al. (2022) considers the community condition of a node with a tailored GCN. Wang et al. (2021) effectively combines the outliers detection ability of one-class SVM and the representation capacity of GNN. Zhou et al. (2021) takes full advantage of the deviation of nodes from the background community to induce the normal node closer to the center of the hypersphere. Moreover, Liu et al. (2021) first adopted the Contrastive approach to anomaly detection on the attributed network. And many schemes based on CoLA are proposed. Jin et al. (2021) deploy both node-node and node-subgraph contrast pairs in the module. Zheng et al. (2021) and Zhang et al. (2022) simultaneously involve information reconstruction and contrastive learning in the training phase and score inference phase.
@@ 2.2 Graph Contrastive Learning
Given a graph, contrastive learning aims to learn networks that map representation in positive pair instances similar to each other, and that embeddings in positive pair instances disagree with each other. Based on the graph transformation type applied in the contrastive learning, the existing work can be divided into three main categories, namely feature, structure and sampling-based transformation. You et al. (2020) systematically studies the impact of many transformation methods, including node attribute masking, random edge perturbation, random node dropping, and random walk-based subgraph sampling. Hassani and Khasahmadi (2020) proves the effectiveness of GDN (Graph diffusion networks) as a view augmentation method. Veličković et al. (2019), Sun et al. (2020), and Hassani and Khasahmadi (2020) validate adequately the Ego-nets sampling as a favorable sample method. By contrast, Qiu et al. (2020) considered that random walk sampling is a stronger method. Most CL-based anomaly detection frameworks employ one or more above-mentioned strategies such that the complicated patterns can be fully modeled.
@@ 4 Methodology
In this section, we introduce the proposed method in detail. The overall pipeline and algorithm are illustrated in Figure 2 and Algorithm 1. SCALA is mainly consists of four components. To be specific, it first calculates the similarity matrix of the node attribute, which is efficient in capturing the relationship of the whole network and the sparsification is conducted to obtain the spar-view (abbreviation for sparse-view), while the origin network is called the dense-view. After that, SCALA sample the subgraph around the target node for both view. Then, the target node and its corresponding subgraph are fed into contrastive modules of two views respectively to gain the discriminative scores. With the attention to easing the negative impact on the quality of graph-level embedding, different readout strategies are operated in two views. After training, SCALA finally integrates scores of the sparsification and the contrastive learning module to gain a final anomaly score.
@@ 4.1 Graph Augmentation via graph sparsification and Subgraph Sampling
Contrastive Learning methods have proved to be of great use in anomaly detection Liu et al. (2021). Graph sparsification is conducted to gain a new view due to such a method can improve the robustness of embeddings for it can prevent the influence from the suspicious relationship. In order to acquire the node-subgraph pair for contrastive learning, a random walk based sampling method is adopted, which is proven to be effective You et al. (2020).
@@ 4.1.1 Graph Sparsification
As mentioned in the introduction section, a node connected by more abnormal edges is more likely to be an anomaly, which provides a new way of qualifying the degree of abnormality. Based on such an idea, we propose a new score calculation method by the process of sparsification, which can be formulated as:
@@ 4.1.2 Random Walk with Restart
Output: Anomaly score mapping $f$($\cdot$).
@@
15:     Calculate the anomaly score $f$($v_{i}$) via Eq.(19).
@@ 4.2 Contrastive Learning Module
In this section, we introduce the contrastive module of SCALA. The contrastive module contains two views. employ The node-graph contrast method is employed on both views which aims to calculate the discriminative score. The main difference between the two views is that different readout strategies are adopted. Furthermore, anonymization is conducted on the attribute of the node in the subgraph corresponding to the target node to prevent target node information in the subgraph from influencing the discrimination.
@@ 4.2.1 Dense-View Contrastive Learning
where $\mathbf{W}_{d}\in\mathbb{R}^{d\times d}$ is the parameter of the discriminator and the $\sigma\left(\cdot\right)$ is the sigmoid activation function that maps the score into range $\left[0,1\right]$. Following the CoLA, the discriminative score of positive pair $s_{i}^{d}$ is supposed to be $1$, by contrast, the discriminative score of negative pair $\tilde{s}_{i}^{d}$ is close to 0. Thus, the BCE (Binary Cross Entropy) loss function is employed to train the module, the loss function of dense-view can be :
@@ 4.2.2 Spar-View Contrastive Learning
where ${\tilde{\mathbf{D}}_{i}^{*}}$ is the degree matrix of $\tilde{\mathbf{A}}_{i}^{*}=\mathbf{A}^{spar}+\mathbf{I}_{P}$. To minimize the influence of graph-level embedding by noise information, we no longer use the unselected Average Readout and propose a simple yet effective similarity-based attention mechanism in spar-view, which can adaptively select useful embedding of the subgraph. Intuitively, if the node in the subgraph is more similar to the target node, the more weight it should have in the process of pooling and vice versa. In order to allow the module to adaptively utilize the similarity information, an MLP is leveraged to apply a homogeneous dimensional mapping to similarity vector $\mathbf{s}_{i}$, which can be formulated as:
@@
As with dense-view, the BCE is employed as our loss function:
@@ 4.3 Anomaly Score Inference
The anomaly scores of the two modules are then unified by the coefficient $\gamma$:
$score_{r}^{con}\left(v_{i}\right)=\left(1-\gamma\right)score_{r}^{con-d}+\gamma score_{r}^{con-s}$ (18)
Notice that it is difficult to adequately capture the context with respect to the target node with only one conduction of sampling. Therefore, multiple samples were taken during the anomaly score inference process and the average of these scores was used as the final score for the whole comparison module:
@@ 5 Experiment
In this section, we conduct extensive experiments on five graph benchmark datasets that are widely used in anomaly detection on attributed networks to verify the performance of SCALA.
@@ 5.1.1 Datasets
To comprehensively evaluate the proposed model, we chose five benchmark datasets including Cora, Citeseer, PubMed Sen et al. (2008), ACM, and DBLP Tang et al. (2008). Due to the shortage of ground truth anomalies in these datasets, we refer to the perturbation scheme following DOMINANT to inject structural anomalies and attribute anomalies for each dataset Skillicorn (2007); Song et al. (2007). The statistics of the datasets are shown in Table 1.
@@ 5.1.2 Baselines and Evaluation Metrics
In this subsection, we compare SCALA with ten well-known baseline methods. The first two models(Radar and ANOMALOUSLi et al. (2017); Peng et al. (2018)) are shallow algorithms, and the rest are based on graph neural networks, including a generative method (DOMINANTDing et al. (2019)), five contrastive methods (CoLA, ANEMONE, GRADATE, NLGAD, and ARISELiu et al. (2021); Jin et al. (2021); Duan et al. (2023a, c, b)), and two hybrid methods that combine generative and contrastive learning (SL-GAD and Sub-CRZheng et al. (2021); Zhang et al. (2022)). ROC-AUC is utilized to measure the performance which is a widely-used anomaly detection metric.
@@ 5.2 Result Analysis
In this subsection, we compare SCALA with ten baseline methods. Figure 3 demonstrates the ROC curves for nine models in five datasets. The AUC of SCALA and baselines are summarized in Table 2. For the results, we have the following conclusions.
We can intuitively find that our proposed model is superior to its competitors on these five datasets. To be specific, SCALA achieves notable AUC gains of 2.66%, 2.64%, 1.83%, 2.25%, and 5.71% on Cora, Citeseer, PubMed, ACM, and DBLP, respectively. As shown in Figure 3, the AUC of SCALA is significantly larger than others.
Non-deep learning-based methods perform worse than other models due to their limited capability to discriminate anomalies from graphs with high-dimensional features and complex structures. The DOMINANT also does not show competitive performance because it aims to recover attribute and structural information rather than directly capture the anomaly information. Compared with previous CL-based methods, CoLA, ANEMONE, GRADATE, NLGAD, and ARISE, SCALA has significant improvements on five datasets. It indicates that the sparsification and the attention mechanism can effectively improve the quality of graph-level embedding. Compared with two hybrid methods that combine reconstruction-based and contrastive-based modules(Sub-CR and SL-GAD), SCALA performs better in anomaly detection. The support of anomaly measurement by sparsification is stronger than attribute reconstruction based on neighbors.
@@ 5.4 Sensitivity Analysis
Figure 4 shows the effect of different $\gamma$ and $\lambda$ values on the model performance. We perform a grid search with stride=0.1 for $\gamma$ and $\lambda$. In practice, we set $\gamma$ to 0.9, 0.8, 0.1, 0.1, 0.6 on Cora, Citeseer, PubMed, ACM, and DBLP. Meanwhile, we set $\lambda$ to 0.1 for DBLP, and the remaining four datasets are all set to 0.2.

### 08 Jan 2024 v2

@@ Abstract
Anomaly detection on attributed networks aims to find the nodes whose behaviors are significantly different from other majority nodes. Generally, network data contains information about relationships between entities, and the condition of a relationship can provide important hints for finding node anomalies. Therefore, how to comprehensively model complex interaction patterns in networks is still a major focus. It can be observed that anomalies in networks violate the homophily assumption. However, most existing studies only consider this phenomenon obliquely rather than explicitly. Besides, the node representation of normal entities can be perturbed easily by the noise relationships introduced by anomalous nodes. To address the above issues, we present a novel contrastive learning framework for anomaly detection on attributed networks, SCALA, aiming to improve the embedding quality of the network and provide a new measurement of qualifying the anomaly score for each node by introducing sparsification into the conventional contrastive method. Extensive experiments are conducted on five real-world datasets and the results show that SCALA consistently outperforms all baseline methods significantly.
@@ 1 Introduction
Tasks on networks have drawn much attention in recent years Tan et al. (2019), particularly for anomaly detection on the network, because there are many complex interaction patterns between entities hard to process Kim et al. (2022). Unlike in other domains, the anomaly detection in networks requires the identification of multiple anomaly types, including attribute anomaly and structure anomaly Ding et al. (2019). The former describes the node with unusual condition of attributes and the latter means the node that has the abnormal situation of the network topology. How to simultaneously identify these anomalies remains a major current focus.
@@
Thanks to the increasing development of GNN which provides a tool to extract the topology and attribute knowledge at the same time Wu et al. (2021). Many researchers introduced GNN into anomaly detection on attributed networks Kim et al. (2022). In general, there are three main categories of GNN methods. First, the Auto-Encoder based method assumes that the attribute and the topology pattern of a normal node can be correctly reconstructed by the decoder while the abnormal nodes cannot Ding et al. (2019). Second, the One-Class SVM based method believes that all the normal nodes can be mapped in the boundary of a hypersphere, while the abnormal node will be expelled out of the boundary Wang et al. (2021). At last, the Contrastive Learning based (CL-based) method determines the degree of node anomaly by calculating the agreement score between the target node and its surrounding neighbor nodes. For normal nodes, their dense embeddings should maintain a matching condition with their positive subgraph embeddings, while abnormal nodes do not retain such a condition Liu et al. (2021). A number of studies have shown that the CL-based method is currently the best method of all, because it can comprehensively and intuitively leverage the matching relationship between the target node and its neighbors Liu et al. (2021); Jin et al. (2021); Zheng et al. (2021); Zhang et al. (2022); Duan et al. (2023a).
@@
It is generally accepted that most nodes of the real-world network satisfy the homophily assumption, which means that connected nodes tend to have a high degree of similarity McPherson et al. (2001). However, as shown in Figure 1(a), the abnormal nodes don’t. Surprisingly, using this discipline as the criteria for spotting anomalies has remained underexplored in previous studies. Besides, as illustrated in Figure 1(b), the noise information of abnormal nodes will cause divergence when obtaining the embedding of the subgraph because most of the CL-based methods inevitably pick the abnormal node during sampling the contrastive pair. Therefore, the agreement calculation will have an under-best performance.
@@
  • We creatively propose a new way to identify anomalies which complement the structural information neglected by the conventional CL-based method.
  • Extensive experiments are conducted on five real-world datasets, whose results show that SCALA consistently outperforms all baseline methods significantly.
@@ 2.1 Anomaly Detection on Attributed Network
With the rapid development of deep learning, researchers also tried to adopt deep learning techniques to detect anomalies in networks. For instance, Ding et al. (2019) adopts an autoencoder to spot anomalies by measuring the reconstruction errors of nodes. Yuan et al. (2021) explicitly introduces the high-order motif structure information into the autoencoder scheme. Luo et al. (2022) considers the community condition of a node with a tailored GCN. Wang et al. (2021) effectively combines the outliers detection ability of one-class SVM and the representation capacity of GCN. Zhou et al. (2021) takes full advantage of the deviation of nodes from the background community to induce the normal node embeddings. Moreover, Liu et al. (2021) first adopted the contrastive approach to anomaly detection on the attributed networks. And many schemes based on CoLA are proposed. Jin et al. (2021) deploy both node-node and node-subgraph contrast pairs in the module. Duan et al. (2023a) investigate the efficiency of subgraph-subgraph contrast pairs. Zheng et al. (2021) and Zhang et al. (2022) simultaneously involve information reconstruction and contrastive learning in the training phase and score inference phase. Duan et al. (2023b) combine the substructure anomaly discerning and the contrastive method. Duan et al. (2023c) take the reliability of the normal node into consideration.
@@ 2.2 Graph Contrastive Learning
Contrastive learning aims to learn networks that map representation in positive pair instances similar to each other, and that embeddings in positive pair instances disagree with each other. You et al. (2020); ssl systematically studies the impact of many transformation methods applied in the contrastive learning, including node attribute masking, random edge perturbation, random node dropping, and random walk-based subgraph sampling. Hassani and Khasahmadi (2020) proves the effectiveness of GDN (Graph diffusion networks) as a view augmentation method. Veličković et al. (2019), Sun et al. (2020), and Hassani and Khasahmadi (2020) validate adequately the Ego-nets sampling as a favorable sample method. By contrast, Qiu et al. (2020) considered that random walk sampling is a stronger method. Most CL-based anomaly detection frameworks employ one or more above-mentioned strategies such that the complicated patterns can be fully modeled.
@@ 4 Methodology
In this section, we introduce the proposed method in detail. The overall pipeline and algorithm are illustrated in Figure 2 and Algorithm 1. SCALA is mainly consists of four components. To be specific, it first calculates the similarity matrix of the node attribute, which is efficient in capturing the relationship of the whole network and the sparsification is conducted to obtain the spar-view (abbreviation for sparse-view), while the origin network is called the dense-view. After that, SCALA sample the subgraph around the target node for both view. Then, the target node and its corresponding subgraph are fed into contrastive modules of two views respectively to calculate the discriminative scores. With the attention to ease the negative impact on the quality of graph-level embedding, different readout strategies are operated in two views. After training, SCALA finally integrates scores of the sparsification and the contrastive learning module to gain a final anomaly score.
@@ 4.1 Graph Augmentation via graph sparsification and Subgraph Sampling
In our approach, the graph sparsification is conducted to obtain a new view due to such a method can improve the robustness of embeddings. In order to acquire the node-subgraph pair for contrastive learning, a random walk based sampling method is adopted, which is proven to be effective You et al. (2020).
@@ 4.1.1 Graph Sparsification
As mentioned in the introduction section, a node connected by more abnormal edges is more likely to be an anomaly, which provides a new way of qualifying the degree of abnormality. Based on such idea, we propose a new score calculation method by the process of sparsification, which can be formulated as:
@@ 4.1.2 Random Walk with Restart
Output: Anomaly score mapping function:$score$($\cdot$).
@@
15:     Calculate the anomaly score $score$($v_{i}$) via Eq.(16)-(19).
@@ 4.2 Contrastive Learning Module
In this section, we introduce the contrastive module of SCALA. The contrastive module contains two views. The node-graph contrast method is employed on both views which aims to calculate the discriminative score. The main difference between the two views is that different readout strategies are adopted. Furthermore, in order to prevent the influence on the discrimination from target node information in the subgraph, the anonymization is conducted on the attribute of the node in the subgraph corresponding to the target node.
@@ 4.2.1 Dense-View Contrastive Learning
where $\mathbf{W}_{d}\in\mathbb{R}^{d\times d}$ is the parameter of the discriminator and the $\sigma\left(\cdot\right)$ is the sigmoid activation function that maps the score into range $\left[0,1\right]$. Following the CoLA, the discriminative score of positive pair $s_{i}^{d}$ is supposed to close to $1$, by contrast, the discriminative score of negative pair $\tilde{s}_{i}^{d}$ is supposed to close to 00. Thus, the BCE (Binary Cross Entropy) loss function is employed to train the module, the loss function of dense-view can be :
@@ 4.2.2 Spar-View Contrastive Learning
where ${\tilde{\mathbf{D}}_{i}^{*}}$ is the degree matrix of $\tilde{\mathbf{A}}_{i}^{*}=\mathbf{A}^{spar}+\mathbf{I}_{P}$. To minimize the influence on graph-level embedding by noise information, we no longer use the unselected Average Readout and propose a simple yet effective similarity-based attention mechanism in spar-view, which can adaptively select useful embedding of the subgraph. Intuitively, if the node in the subgraph is more similar to the target node, the more weight it should have in the process of pooling and vice versa. In order to allow the module to adaptively utilize the similarity information, an MLP is leveraged to apply a homogeneous dimensional mapping to similarity vector $\mathbf{s}_{i}$, which can be formulated as:
@@
As with dense-view, the BCE is employed as the loss function:
@@ 4.3 Anomaly Score Inference
The anomaly scores of the two views are then unified by the coefficient $\gamma$:
$score^{con}\left(v_{i}\right)=\left(1-\gamma\right)score^{con-d}+\gamma score^{con-s}$ (18)
Notice that it is difficult to adequately capture the context with respect to the target node with only one conduction of sampling. Therefore, multiple samples were taken during the anomaly score inference process and the average of these scores was used as the final score for the whole contrastive module:
@@ 5 Experiment
In this section, we conduct extensive experiments on five network benchmark datasets that are widely used in anomaly detection on attributed networks to verify the performance of SCALA.
@@ 5.1.1 Datasets
To comprehensively evaluate the proposed model, we choose five benchmark datasets including Cora, Citeseer, PubMed Sen et al. (2008), ACM, and DBLP Tang et al. (2008). Due to the shortage of ground truth anomalies in these datasets, we refer to the perturbation scheme following DOMINANT to inject structural anomalies and attribute anomalies for each dataset Ding et al. (2019); Skillicorn (2007); Song et al. (2007). The statistics of the datasets are shown in Table 1.
@@ 5.1.2 Baselines and Evaluation Metrics
In this subsection, we compare SCALA with ten well-known baseline methods. The first two models (Radar and ANOMALOUS Li et al. (2017); Peng et al. (2018)) are shallow algorithms, and the rest are based on graph neural networks, including a Auto-Encoder based method (DOMINANT Ding et al. (2019)), five CL-based methods (CoLA, ANEMONE, GRADATE, NLGAD, and ARISE Liu et al. (2021); Jin et al. (2021); Duan et al. (2023a, c, b)), and two hybrid methods that combine Auto-Encoder and CL based learning (SL-GAD and Sub-CR Zheng et al. (2021); Zhang et al. (2022)). ROC-AUC is utilized to measure the performance which is a widely-used anomaly detection metric.
@@ 5.2 Result Analysis
In this subsection, we compare SCALA with ten baseline methods. Figure 3 demonstrates the ROC curves for 11 models in five datasets. The AUC of SCALA and baselines are summarized in Table 2. For the results, we have the following conclusions.
We can intuitively find that our proposed model is superior to its competitors on these five datasets. To be specific, SCALA achieves notable AUC gains of 2.66%, 2.64%, 1.60%, 2.25%, and 5.71% on Cora, Citeseer, PubMed, ACM, and DBLP, respectively. As shown in Figure 3, the AUC of SCALA is significantly larger than others.
Shallow methods perform worse than other models due to their limited capability to discriminate anomalies from graphs with high-dimensional features and complex structures. The DOMINANT also does not show competitive performance because it aims to recover attribute and structural information rather than directly capture the anomaly information. Compared with previous CL-based methods, CoLA, ANEMONE, GRADATE, NLGAD, and ARISE, SCALA has significant improvements on five datasets. It indicates that the sparsification and the attention mechanism can effectively improve the quality of graph-level embedding. Compared with two hybrid methods that combine reconstruction-based and contrastive-based modules (Sub-CR and SL-GAD), SCALA performs better in anomaly detection. This shows that the support of anomaly measurement by sparsification is stronger than attribute reconstruction based on neighbors.
@@ 5.4 Sensitivity Analysis
Figure 4 shows the effect of different $\gamma$ and $\lambda$ values on the model performance. We perform a grid search with stride 0.1 for $\gamma$ and $\lambda$. In practice, we set $\gamma$ to 0.9, 0.8, 0.1, 0.1, 0.6 on Cora, Citeseer, PubMed, ACM, and DBLP. Meanwhile, we set $\lambda$ to 0.1 for DBLP, and the remaining four datasets are all set to 0.2.